News by Date
News by Category
Solar System
White Dwarfs
Neutron Stars
Black Holes
Milky Way Galaxy
Normal Galaxies
Galaxy Clusters
Cosmology/Deep Field
Press Resources
Status Reports
Press Advisories
Image Releases
Release Guidelines
Image Use Policy
Web Shortcuts
Chandra Blog
RSS Feed
Email Newsletter
News & Noteworthy
Image Use Policy
Questions & Answers
Glossary of Terms
Download Guide
Get Adobe Reader
Chandra @ NASA
Visit the Chandra pages at the NASA portal (opens in new window)
Image Use
Image Use Policy & Request Form
Guidelines for utilizing images, applets, movies, and animations featured in this Web Site.
Operations CXO Status Report

Tuesday 5 October 1999 9.00am EST

During the last 24 hours Chandra completed a series of calibration measurements of AR Lac with the HRC, and completed a pre-radiation belt ACIS CTI measurement. The HRC door was partially closed in preparation for belt passage. The spacecraft entered the belts at 7.30AM 10/5 and will enter the 8th eclipse of the season at 11:15AM 10/5. The eclipse will last 58 minutes.

Following belt exit at 6.45PM 10/5 an ACIS CTI measurement will be conducted then a long (145 Ks) observation of SN1987A with ACIS-S and the High Energy Transmission Grating. This GTO observation will run throughout the next orbit.

The spacecraft systems continued to operate nominally.

Return to Status Reports